OpenStudy (anonymous):

Can someone guide me solving this system of equation? y = x + 3 2x + y = 9

OpenStudy (ajanijones):

What do you need help with?

OpenStudy (anonymous):

...Solving the system of equation.

OpenStudy (anonymous):

@agent0smith @mathstudent55 @hartnn @ganeshie8 @kelliegirl33 @CGGURUMANJUNATH @radar

OpenStudy (cggurumanjunath):

PUT Y=X+3 IN 2X+Y=9

OpenStudy (cggurumanjunath):

2X+(X+3)=9

OpenStudy (cggurumanjunath):

NOW CAN U SOLVE FOR X ? @Darry&Gator

OpenStudy (cggurumanjunath):

@Darry&Gator

OpenStudy (cggurumanjunath):

IS IT CLEAR OR NOT ?

OpenStudy (anonymous):

Um, I think I can...do i distribute 2x to x and 3?

OpenStudy (cggurumanjunath):